OpenStudy (elizabeth):

what is .00000001 as a fraction

OpenStudy (mathteacher1729):

\[.1 =\frac{1}{10}\] \[.01 =\frac{1}{100}\] \[.001 =\frac{1}{1000}\] \[.0001 =\frac{1}{10000}\] etc...

OpenStudy (anonymous):

1/100000000 count the number of zeros before the decimal and first nonzero digit. Your fraction should be 1 over a power of 10 that has one more zero

OpenStudy (elizabeth):

thanks...